Sound with intensity larger than 120 dB appears painful to a person. A small speaker delivers 2.0 W of audio output. How close can the person get to the speaker without hurting his ears?

Grade:11

1 Answers

Aditi Chauhan
9 years ago
Sol. If sound level = 120 dB, then I = intensity = 1 W/m^2 Given that, audio output = 2W Let the closest distance be x. So, intensity = (2 / 4πx^2) = 1 ⇒ x^2 = (2/2π) ⇒ x = 0.4 m = 40 cm.